OpenStudy (anonymous):

(-2x+6)^1/5=(-8+10x)^1/5

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

anyone know how to do this?

OpenStudy (anonymous):

Since they are bot to the 1/5 power you can eliminate that and solve for (-2x+6) = (-8+10x) Do you need help with that?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

do you use foil?

OpenStudy (anonymous):

-2x +6 = -8 +10x. We need to get the variables ( x values) to one side of the equation and the constants to the other side. So how would you do that?

OpenStudy (anonymous):

foil is for multiplying

OpenStudy (anonymous):

okay so you just add 2 to each side and add 8 to the 6

OpenStudy (anonymous):

Yes and solve for x. What do you get?

OpenStudy (anonymous):

so then i get 14+12x?

OpenStudy (anonymous):

and then divide 12 to each side?

OpenStudy (anonymous):

i mean 14=12x

OpenStudy (anonymous):

yes and divide both sides by 12

OpenStudy (anonymous):

okay but why do you just forget the 1/5?

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

Sorry had to take care of something. In equations that have the same exponent they would cancel out. It is a rule.

OpenStudy (anonymous):

oh okay thanks!